OpenStudy (anonymous):

What type of polynomial is x(x - 4)(x - 5)?

OpenStudy (anonymous):

I think it is a cubic(3rd degree) binomial

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@mathslover @sleepyjess

OpenStudy (anonymous):

@alekos

OpenStudy (alekos):

Cubic trinomial

OpenStudy (solomonzelman):

Yes, you are correct, it is a cubic polynomial (3rd degree polynomial).

OpenStudy (anonymous):

Should I just say polynomial or should I say trinomial? Is it a trinomial? I thought it was a binomial.

OpenStudy (solomonzelman):

trinomial = is 3 terms. how many terms will you have though?

OpenStudy (anonymous):

x(x - 4)(x - 5) i know for sure that (x - 4) is a term and that (x - 5) is a term but I was't sure about that x in front

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle x(x - 4)(x - 5)= }\) \(\large\color{black}{ \displaystyle x(x ^2 -4x-5x+20)= }\) \(\large\color{black}{ \displaystyle x(x ^2 -9x+20)= }\) \(\large\color{black}{ \displaystyle x ^3 -9x^2+20x }\)

OpenStudy (solomonzelman):

yes, 3 terms :) ... any questions?

OpenStudy (anonymous):

ahh okay I see what you did there. Thank you very much :)

OpenStudy (solomonzelman):

polynomial with a degree of 3. --> Cubic 3 terms in the expression. --> Trinomial

OpenStudy (solomonzelman):

cubic trinomial... yw